OpenStudy (anonymous):

-2<2x -4 <4 How do I graph this on a number line? By the way, the < has a line under it.

OpenStudy (solomonzelman):

add 4 to each part divide each part by 2

OpenStudy (anonymous):

Do you mean negative two?

OpenStudy (solomonzelman):

no, I mean just by 2.

OpenStudy (solomonzelman):

just follow the steps I suggested (please)

OpenStudy (anonymous):

So 1<x<4

OpenStudy (solomonzelman):

yes, there we go !

OpenStudy (anonymous):

Do I subtract the 1?

OpenStudy (solomonzelman):

No, you got the answer, you just need to graph it. When you graph it on the number line, you will draw 2 open dots, one at 1, and the other at 4. (open because 1 and 4 are not part of the solution, right?) And then connect the 2 dots.

OpenStudy (anonymous):

Right, okay, I have another problem like this, will you stay in this so I can work it out with you?

OpenStudy (solomonzelman):

yes, sure

OpenStudy (anonymous):

6x - 4 < -34 or 3x +10 > 10 Why does it say or?!

OpenStudy (anonymous):

As in, I pick one to work out? Or do both?

OpenStudy (solomonzelman):

it probably means that it is just in one question.

OpenStudy (solomonzelman):

work both, one by one.

OpenStudy (solomonzelman):

Let start from \(\normalsize\color{blue}{ 6x - 4 < -34\LARGE\color{white}{ \rm │ }}\) .

OpenStudy (anonymous):

http://gyazo.com/ba7aa9b534fbe4742fc798fd6e94288b here is a screencap of it btw, and okay

OpenStudy (solomonzelman):

okay, so can you just solve the \(\normalsize\color{black}{ 6x - 4 < -34\LARGE\color{white}{ \rm │ }}\) ?

OpenStudy (anonymous):

so now its 6x<-30

OpenStudy (solomonzelman):

It says `or` (by the way) because the 2 equations are one graph on the number line. That is correct for the first step, now.. divide both sides by 6.

OpenStudy (anonymous):

So x < 5

OpenStudy (anonymous):

So is that the solution?

OpenStudy (solomonzelman):

no close, x<-5

OpenStudy (solomonzelman):

see why ?

OpenStudy (anonymous):

ah, my mistake. Because the 30 was negative.

OpenStudy (solomonzelman):

yes, right... and we get x<-5 Now, can you please work the \(\normalsize\color{blue}{ 3x +10 > 10 \LARGE\color{white}{ \rm │ }}\) ?

OpenStudy (anonymous):

no solution?

OpenStudy (solomonzelman):

X>0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to do absolute value equations?

OpenStudy (solomonzelman):

any x-value that is more than zero will make `3x` (also) be more than zero, and thus the right side will be more than 10. But if x is a zero, then you get 10>10 right? and this is FALSE and certainly if x is less than a zero i.e. a negative number.

OpenStudy (solomonzelman):

Yes.. hopefully :)

OpenStudy (anonymous):

http://gyazo.com/263afbec7b48b064737015c5e58c72e7

OpenStudy (anonymous):

or would you rather me type it

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2│3x-5│-8=8\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2│3x-5│-8=8\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 2│3x-5│-8\color{magenta}{+8}=8\color{magenta}{+8}\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 2│3x-5│=16\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 2│3x-5│\color{green}{\div 2}=16\color{green}{\div 2}\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ │3x-5│=8\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 3x-5=8~~~~~~~~~or~~~~~~~3x-5=-8\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

you can do it from here.

OpenStudy (anonymous):

Yeah I did that then I got 3x-5=8 add the 5, so 3x=13 but I dont know what to do with that because you cant divide 13 by 3.

OpenStudy (solomonzelman):

you CAN divide 13 by 3, and you can write the answer either as a fraction, or as a decimal. I recommend to do it as a fraction, so that it will be exact.

OpenStudy (anonymous):

is it 4 1/3 or 4.33?

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

So x=13/3 or x=4 1/13 (improper fraction) (mixed number)